A train of mass 200kN has a frictional resistance of 5N per kN. The speed of the


train, at the top of an inclined of 1 in 80 is 45 km/h. Find the speed of the train after


running down the incline for 1km.

1
Expert's answer
2022-02-10T16:15:09-0500

Given:

"W=200\:\rm kN"

"F_f=1000\:\rm N"

"\sin\theta=1/80"

"l=1000\:\rm m"


The net force

"F=W\sin\theta-F_f\\
=200\:000*1/80-1000=1500\:\rm N"

The acceleration of a train

"a=\frac{F}{m}=\frac{Fg}{W}=\frac{1500*9.8}{200\: 000}=0.0735\:\rm m/s^2"

The final velocity

"v=\sqrt{2la}=\sqrt{2*1000*0.0735}=12.1\:\rm m/s"